OpenStudy (anonymous):

4x – 5y = 9 I need help rememberin how to do this

OpenStudy (anonymous):

I am solvin for y

OpenStudy (studygurl14):

just rearrange the equation nd isolate the y. FIRST STEP: 4x - 5y = 9 -5y = 9 - 4x See? Just keep going.

OpenStudy (opcode):

\[4x – 5y = 9\] Obtain \(y\) alone: \[4x = 9 + 5y\] \[4x - 9 = 5y\] Divide both sides by \(5\): \[4x - 9 = 5y \Rightarrow y = \dfrac{4x - 9}{5}\]

OpenStudy (anonymous):

oh thats right! Thank you both bunches!

OpenStudy (anonymous):

@StudyGurl14 will you give @Opcode a medal for helpin too?

OpenStudy (studygurl14):

sure

OpenStudy (anonymous):

thank you